OpenStudy (anonymous):

Solve for the variable: -5g – 6 = 19 A. -7 B. -5 C. 5 D. 7

OpenStudy (anonymous):

-5g-6=19 -5g=19+6 -5g=25 -g=5 g=-5

OpenStudy (anonymous):

what did you do exactly?

OpenStudy (anonymous):

ok so anytime you have an "equation" there are 2 sides...

OpenStudy (anonymous):

let me start with a more basic example

OpenStudy (anonymous):

x+2x=1+2

OpenStudy (anonymous):

there is a left and right side

OpenStudy (anonymous):

you combine like terms

OpenStudy (anonymous):

x+2x=1+2, 3x=3, does this make sense so far

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

could you write it out?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

here is an equation: x+2x=1+2 step 1: x+2x=1+2 step 2: 3x=3 step 3: x=3/3 step 4: x=1. Answer: x = 1

OpenStudy (anonymous):

okay i am starting to get the idea

OpenStudy (anonymous):

now lets try the problem above...

OpenStudy (anonymous):

-5g-6=19.....

OpenStudy (anonymous):

Equation: Let -5g-6=19, Step 1: -5g-6=19, Step 2: -5g=19+6... anytime you want to move a number to the other side, the sign becomes the opposite sign it was on the previous side, in other words, you change the sign... Step 3: -5g=25, Step 4: -5g/5=25/5..... I divided both sides by 5, whatever you do to one side, you must do to the other side..... Step 5: -g=5 Step 6: g=-5 Hence the answer is -5.